work and return back. You may attempt to take the midterm twice, however you est and you cannot save your grade will be the average of the two if you do so! There are 16 multiple choice questions, 5 pts. each. Then you will have to answer two free response questions. type as many details as possible, so that I maybe able to give partial credit. Indicate your final answer as The ans is: for each section. Thank you. Good Luck D | Question 2 5 pts The distance of the Earth from the sun is 93 000 000 miles. If there are 3.15 × 107 s in one year, find the speed of the Earth in its orbit about the sun. O 1674 miles/min O 2232 miles/min O 1116 miles/min 0 557 miles/min 4 Previous Next
Answer #1

Circumferential distance travelled in one year = 2 * pi* 93000000 miles

D = 584336233.5677 miles

Speed of Earth around the sun it its orbit = D/t = 584336233.5677 miles/3.15 *10^7 miles/s

Speed = 18.55 miles/second

Speed = 18.55 miles/second *60 s/1 min = 1116 miles/min

Answer is 1116 miles/min